# UrbanNav: Learning Language-Guided Urban Navigation from Web-Scale Human Trajectories

### v1

@@ Abstract
Navigating complex urban environments using natural language instructions poses significant challenges for embodied agents, including noisy language instructions, ambiguous spatial references, diverse landmarks, and dynamic street scenes. Current visual navigation methods are typically limited to simulated or off-street environments, and often rely on precise goal formats, such as specific coordinates or images. This limits their effectiveness for autonomous agents like last-mile delivery robots navigating unfamiliar cities. To address these limitations, we introduce UrbanNav, a scalable framework that trains embodied agents to follow free-form language instructions in diverse urban settings. Leveraging web-scale city walking videos, we develop an scalable annotation pipeline that aligns human navigation trajectories with language instructions grounded in real-world landmarks. UrbanNav encompasses over 1,500 hours of navigation data and 3 million instruction-trajectory-landmark triplets, capturing a wide range of urban scenarios. Our model learns robust navigation policies to tackle complex urban scenarios, demonstrating superior spatial reasoning, robustness to noisy instructions, and generalization to unseen urban settings. Experimental results show that UrbanNav significantly outperforms existing methods, highlighting the potential of large-scale web video data to enable language-guided, real-world urban navigation for embodied agents.
@@ 1 Introduction
Language-guided navigation in real-world urban environments is a cornerstone capability for autonomous agents, enabling applications such as last-mile delivery robots, autonomous vehicles, and assistive robotics. Unlike structured indoor spaces or synthetic simulation settings, urban scenes are inherently dynamic, featuring diverse terrains, unpredictable obstacles, and dense pedestrian interactions (shah2023gnm). To operate effectively in such environments, embodied agents must not only reason about spatial layouts and adhere to implicit social norms but also interpret ambiguous human instructions, such as “go to the cafe by the old bridge” or “move to the bookstore opposite the park.” These often vague or context-specific directives require sophisticated reasoning, making language-guided urban navigation a complex yet essential task for deploying autonomous agents in real-world cities (gao2025openfly).
@@
Prior research (shah2023vint; sridhar2024nomad; yu2025cnavselfevolvingcontinualobject) on visual navigation has made substantial progress in simulation and indoor domains. Classical methods (muhlbauer2009navigation; kummerle2013navigation) combine SLAM and modular planning to achieve goal-oriented navigation, while more recent works (ehsani2024spoc; zeng2024poliformer) leverage reinforcement learning or imitation learning within high-fidelity simulators. These advances have enabled impressive results in point goal or object goal navigation tasks. However, current approaches are still constrained by their reliance on precise goal specifications, such as GPS coordinates or target images and limited diversity.
@@
The complexity of language-guided urban navigation stems from the need to align natural language instructions with real-world spatial (schumann2022analyzing). In real urban environments, users typically provide free-form directions such as “deliver to the building near the park fountain,” which often reference salient visual landmarks. Human navigation often relies on such landmarks and contextual cues embedded in language, which are challenging to model in simulations or small-scale datasets. Although collecting expert trajectories via teleoperation is an adopted approach, in practice it is constrained by limited data diversity and high annotation costs, hindering generalizability across varied urban scenarios. To overcome these challenges, we propose UrbanNav, a scalable framework illustrated in Fig. LABEL:fig:overview, that leverages web-scale human navigation trajectories from city walking videos to train embodied agents for language-guided urban navigation. UrbanNav answers two key questions associated with learning from unstructured web videos:
@@
First, are all video segments suitable for training embodied agents? Many clips exhibit viewpoint divergence, where camera orientation deviates from the direction of motion—contrary to the forward-facing perspective of robots—or capture unsafe behaviors, such as weaving though dense crowds, posing risks for robotic deployment (bar2025navigation). UrbanNav tackles this by employing a filtering pipeline, leveraging visual odometry to estimate camera pose and detect misalignment between viewpoint and trajectory, and integrating object detector to identify and exclude segments with unsafe interactions. This ensures only high-quality, robot-compatible data informs training.
@@
Second, how to obtain instruction-action supervision from in-the-wild videos for imitation learning? Manual annotation of such videos is infeasible, necessitating an automated, scalable approach (he2025cameractrl). UrbanNav addresses this through a sophisticated pipeline. Using off-the-shelf visual odometry models, we extract egocentric trajectories with pose information from video sequences. Next, a large vision-language model (VLM) detects contextually relevant landmarks in urban scenes, followed by another VLM that produce natural, navigation-oriented instructions grounded in these landmarks. This pipeline yields a dataset of over 1,500 hours of navigation data and 3 million instruction-trajectory-landmark triplets, enabling robust, scalable language-guided navigation policies for diverse urban environments.
@@ 2 Related Works
The research on indoor navigation has greatly advanced with the increasing realism and diversity of simulators (Matterport3D; savva2019habitat; kolve2017ai2; deitke2022). In these environments, some works focus on improving robotic navigation capabilities through imitation learning using expert demonstration data (ehsani2024spoc) or online reinforcement learning (zeng2024poliformer). Other works (zhou2024navgpt; qiao2024open; zhang2025flexvln; ding2025lavira) have attempted to leverage the vast prior knowledge and strong generalization capabilities of LLM for zero-shot navigation reasoning. However, these methods suffer significant performance degradation (gervet2023navigating) and low efficiency (zhu2024minivln; zhang2025cosmo) when deployed on physical robotic platforms.
@@
Directly using real-world navigation data (hirose2018gonet; shah2021rapid; karnan2022socially; hirose2023sacson) to construct training samples for supervised learning is an efficient approach. Some works (shah2023gnm; shah2023vint; shah2022viking) train models on a mix of real-robot navigation datasets, enabling direct deployment on different robotic platforms with strong generalization capabilities, while others (sridhar2024nomad; bar2025navigation; dong2025unifiedworldmodelsmemoryaugmented) employ diffusion models to generate trajectories or leverage the model’s imagination of the future to assist navigation. However, manually collecting such data is costly and lacks diversity. This work automates the construction of navigation datasets using web-scale data to overcome this limitation.
@@
The task of using text descriptions as navigation instructions has been widely studied. Vision-and-Language Navigation (anderson2018vision; qi2020reverie; krantz2020beyond; liu2025groundingmate) requires the agent to follow fine-grained navigation instructions to move within a scene. Object Goal Navigation (chaplot2020object; wani2020multion; cai2024bridging; sadek2023multi; yokoyama2024hm3d) aims to enable agent to efficiently locate specific objects in the environment based on textual descriptions. In contrast, we focus on outdoor scenarios and train the language-guided navigation model using only real-world data, thereby avoiding this gap and enabling the model to learn navigation rules and environmental affordances inherent in urban environments.
@@
Learning from web-scale video data has advanced language and vision tasks (wang2024scaling; he2024omnih2o; he2025asap). However, a key challenge is the lack of action labels for navigation tasks. LeLaN (hirose2024lelan) uses vision-language model prompting and pretrained navigation models to generate action labels, while (liu2025citywalker) applies visual odometry to annotate video frames with pose information. However, LeLaN focuses on close-range object navigation within indoor static scenes, making it incapable of handling complex urban environments, and the constructed dataset does not encompass obstacle avoidance capabilities. In contrast, UrbanNav distinguishes itself by implementing a robust filtering pipeline to select robot-compatible data, ensuring viewpoint alignment and safe navigation behaviors. Additionally, our approach generates language-grounded instructions tied to landmarks and trajectories, enabling scalable, cost-effective, and high-quality action label generation for imitation learning in complex urban navigation tasks.
@@ 3.2 Labeling In-the-Wild Videos
We uniformly segment the raw videos into 2-minute clips, each representing a candidate navigation trajectory. To extract pose information, we employ the state-of-the-art visual odometry model DPVO (teed2023deep). The first frame of each clip is taken as the origin of a local world coordinate system, and we annotate the camera pose of each subsequent frame relative to it. Although visual odometry may suffer from cumulative drift across long horizons, this effect is largely mitigated in our case: the policy is trained to predict future actions based on only the past 8 steps, operating within short temporal windows where VO remains locally consistent. This design, following the practice in (liu2025citywalker), enables reliable supervision from noisy egocentric videos. As a result, we obtained a total of 106,603 egocentric navigation trajectories with pose labels, spanning 3,553 hours.
@@
A crucial yet often overlooked issue in learning from in-the-wild walking videos is that not all segments are suitable for training embodied agents. Robotic platforms typically employ forward-facing, fixed-view cameras, requiring visual inputs that maintain consistent alignment with motion direction. However, Unlike robots, human walkers exhibit flexible body dynamics and frequently adjust their head orientation, resulting in significant pitch variations and misalignment between the camera viewpoint and the actual trajectory. Such discrepancies render many video clips incompatible with robotic requirements. Additionally, some segments feature densely crowded scenes with close pedestrian interactions, which pose safety risks for robotic deployment. Unlike prior work (liu2025citywalker), which often overlooks these issues and risks degrading policy performance, our approach employs a robust filtering pipeline to select predominantly robot-compatible data for training.
First, we estimate the per-frame camera pitch angle and reject trajectories exhibiting excessive vertical oscillations—specifically, those with pitch variation beyond $15^{\circ}$. We further analyze the alignment between movement direction and viewing direction using a sliding window, and discard segments with significant directional divergence (over $60^{\circ}$), which commonly results from abrupt head turns or side glances. Second, to eliminate trajectories captured in overly crowded areas, we apply YOLOv10 (wang2024yolov10) to detect pedestrians. Based on empirical observations, we discard any trajectory where more than five people appear in a single frame and such occurrences happen in more than three frames, indicating sustained dense proximity. After applying these filtering steps, we retain 47,008 high-quality, robot-compatible trajectories spanning 1,566 hours, which serve as the core data for training.
@@
We leverage Qwen2.5-VL-72B (bai2025qwen2), a state-of-the-art VLM, to automatically detect and localize candidate landmarks in video frames via prompted queries aligned with the above criteria. The model outputs both bounding boxes and preliminary landmark names. To ensure quality, we manually review and filter out low-confidence or ambiguous annotations. For the retained landmarks, we prompt Qwen2.5-VL-72B to generate concise and descriptive natural language instructions. Through this process, we obtain a total of 3 million landmark annotations, each paired with a bounding box and a language description. On average, each trajectory contains 65 identified landmarks, with the mean description length being 17 words.
Method
Test Seen Test Unseen
@@ 3.3 Policy Architecture and Training
As illustrated in Fig. 3, our policy model predicts a sequence of future egocentric positions based on both language instructions and visual observations. The input to the model comprises four components: (1) the language instruction, (2) the current visual observation, (3) the past $k$ visual observations. It builds upon previous works (hirose2024lelan; liu2025citywalker). We use CLIP (radford2021learning) to encode the language instruction and DINOv2 (oquab2023dinov2) to extract features from all visual frames. Both encoders are kept frozen during training. To ground the current observation in the instruction, we apply a FiLM module (perez2018film) to modulate the current visual embedding using the language features. All input tokens—including language embeddings, FiLM-modulated current visual features and historical visual observation—are concatenated and fed into a Transformer encoder. This encoder captures both temporal dynamics and cross-modal interactions, and outputs a predicted future trajectory in egocentric coordinates. To enable smoother and more anticipatory navigation, we adopt a multi-step prediction framework following prior work such as (sridhar2024nomad). At each time step, the model predicts the agent’s waypoints for the next $k=8$ steps using a lightweight action head.
@@ 4.1 Experimental Setup
We compare our model against several prominent policies previously for real-world navigation. To ensure a fair comparison, we adapt these baselines to our language-guided task. For instance, NoMaD (sridhar2024nomad) and ViNT (shah2023vint) were originally designed for image-goal navigation, and lacked native support for textual instructions. To adapt them, we augmented their architectures by encoding the natural language instruction with CLIP (radford2021learning), concatenating the resulting text features with the visual features and passing them through a fusion layer, while the core network remained unchanged. Similarly, LeLaN (hirose2024lelan), originally developed for indoor object-goal navigation with textual inputs, predicts robot linear and angular velocities. To align its output space with our task formulation for consistent evaluation, we adapted its output head to directly regress the future navigation waypoints.
@@
Given the challenges of evaluating end-to-end task completion in real-world environments without autoregressive execution, we designed a comprehensive evaluation protocol that includes both offline benchmarking and real-world deployment. (1) Offline Evaluation. For offline evaluation, we assess the model’s performance on a held-out validation set. The model is given a sequence of historical observations, positions, and a natural language instruction. Its task is to predict a future trajectory, and we measure the deviation between this prediction and the ground-truth trajectory. Following prior work (liu2025citywalker), we use the average orientation error (AOE) and maximum average orientation error (MAOE) to measure the directional alignment between predicted and ground-truth actions. To better capture local path-following capabilities essential for obstacle avoidance, we introduce two additional metrics: average distance error (ADE), which calculates the mean L2-distance between predicted and ground-truth positions, and maximum average distance error (MADE), which represents the Fréchet distance (alt1995computing) in a discrete setting. For these metrics, a lower value indicates better performance. A more detailed definition of these metrics is provided in the appendix. (2) Real-World Deployment. To validate the policy’s real-world generalization, we deploy the model on a physical robot platform. The primary metric for this evaluation is the navigation success rate, which measures the percentage of trials where the robot successfully reaches the specified goal landmark without any collisions.

### v2

@@ Abstract
Navigating complex urban environments using natural language instructions poses significant challenges for embodied agents, including noisy language instructions, ambiguous spatial references, diverse landmarks, and dynamic street scenes. Current visual navigation methods are typically limited to simulated or off-street environments, and often rely on precise goal formats, such as specific coordinates or images. This limits their effectiveness for autonomous agents like last-mile delivery robots navigating unfamiliar cities. To address these limitations, we introduce UrbanNav, a scalable framework that trains embodied agents to follow free-form language instructions in diverse urban settings. Leveraging web-scale city walking videos, we develop an scalable annotation pipeline that aligns human navigation trajectories with language instructions grounded in real-world landmarks. UrbanNav encompasses over 1,500 hours of navigation data and 3 million instruction-trajectory-landmark triplets, capturing a wide range of urban scenarios. Our model learns robust navigation policies to tackle complex urban scenarios, demonstrating superior spatial reasoning, robustness to noisy instructions, and generalization to unseen urban settings. Experimental results show that UrbanNav significantly outperforms existing methods, highlighting the potential of large-scale web video data to enable language-guided, real-world urban navigation for embodied agents. All code and data are released at https://github.com/CASIA-IVA-Lab/UrbanNav.
@@ 1 Introduction
Language-guided navigation in real-world urban environments is a cornerstone capability for autonomous agents, enabling applications such as last-mile delivery robots, autonomous vehicles, and assistive robotics. Unlike structured indoor spaces or synthetic simulation settings, urban scenes are inherently dynamic, featuring diverse terrains, unpredictable obstacles, and dense pedestrian interactions (Shah et al. 2023a). To operate effectively in such environments, embodied agents must not only reason about spatial layouts and adhere to implicit social norms but also interpret ambiguous human instructions, such as “go to the cafe by the old bridge” or “move to the bookstore opposite the park.” These often vague or context-specific directives require sophisticated reasoning, making language-guided urban navigation a complex yet essential task for deploying autonomous agents in real-world cities (Gao et al. 2025).
@@
Prior research (Shah et al. 2023b; Sridhar et al. 2024; Yu et al. 2025) on visual navigation has made substantial progress in simulation and indoor domains. Classical methods (Muhlbauer et al. 2009; Kümmerle et al. 2013) combine SLAM and modular planning to achieve goal-oriented navigation, while more recent works (Ehsani et al. 2024; Zeng et al. 2024) leverage reinforcement learning or imitation learning within high-fidelity simulators. These advances have enabled impressive results in point goal or object goal navigation tasks. However, current approaches are still constrained by their reliance on precise goal specifications, such as GPS coordinates or target images and limited diversity.
@@
The complexity of language-guided urban navigation stems from the need to align natural language instructions with real-world spatial (Schumann and Riezler 2022). In real urban environments, users typically provide free-form directions such as “deliver to the building near the park fountain,” which often reference salient visual landmarks. Human navigation often relies on such landmarks and contextual cues embedded in language, which are challenging to model in simulations or small-scale datasets. Although collecting expert trajectories via teleoperation is an adopted approach, in practice it is constrained by limited data diversity and high annotation costs, hindering generalizability across varied urban scenarios. To overcome these challenges, we propose UrbanNav, a scalable framework illustrated in Fig. LABEL:fig:overview, that leverages web-scale human navigation trajectories from city walking videos to train embodied agents for language-guided urban navigation. UrbanNav answers two key questions associated with learning from unstructured web videos:
@@
First, are all video segments suitable for training embodied agents? Many clips exhibit viewpoint divergence, where camera orientation deviates from the direction of motion—contrary to the forward-facing perspective of robots—or capture unsafe behaviors, such as weaving though dense crowds, posing risks for robotic deployment (Bar et al. 2025). UrbanNav tackles this by employing a filtering pipeline, leveraging visual odometry to estimate camera pose and detect misalignment between viewpoint and trajectory, and integrating object detector to identify and exclude segments with unsafe interactions. This ensures only high-quality, robot-compatible data informs training.
@@
Second, how to obtain instruction-action supervision from in-the-wild videos for imitation learning? Manual annotation of such videos is infeasible, necessitating an automated, scalable approach (He et al. 2025a). UrbanNav addresses this through a sophisticated pipeline. Using off-the-shelf visual odometry models, we extract egocentric trajectories with pose information from video sequences. Next, a large vision-language model (VLM) detects contextually relevant landmarks in urban scenes, followed by another VLM that produce natural, navigation-oriented instructions grounded in these landmarks. This pipeline yields a dataset of over 1,500 hours of navigation data and 3 million instruction-trajectory-landmark triplets, enabling robust, scalable language-guided navigation policies for diverse urban environments.
@@ 2 Related Works
The research on indoor navigation has greatly advanced with the increasing realism and diversity of simulators (Chang et al. 2017; Savva et al. 2019; Kolve et al. 2017; Deitke et al. 2022). In these environments, some works focus on improving robotic navigation capabilities through imitation learning using expert demonstration data (Ehsani et al. 2024) or online reinforcement learning (Zeng et al. 2024). Other works (Zhou et al. 2024; Qiao et al. 2024; Zhang et al. 2025a; Ding et al. 2025) have attempted to leverage the vast prior knowledge and strong generalization capabilities of LLM for zero-shot navigation reasoning. However, these methods suffer significant performance degradation (Gervet et al. 2023) and low efficiency (Zhu et al. 2024; Zhang et al. 2025b) when deployed on physical robotic platforms.
@@
Directly using real-world navigation data (Hirose et al. 2018; Shah et al. 2021; Karnan et al. 2022; Hirose et al. 2023) to construct training samples for supervised learning is an efficient approach. Some works (Shah et al. 2023a, b; Shah and Levine 2022) train models on a mix of real-robot navigation datasets, enabling direct deployment on different robotic platforms with strong generalization capabilities, while others (Sridhar et al. 2024; Bar et al. 2025; Dong et al. 2025) employ diffusion models to generate trajectories or leverage the model’s imagination of the future to assist navigation. However, manually collecting such data is costly and lacks diversity. This work automates the construction of navigation datasets using web-scale data to overcome this limitation.
@@
The task of using text descriptions as navigation instructions has been widely studied. Vision-and-Language Navigation (Anderson et al. 2018; Qi et al. 2020; Krantz et al. 2020; Liu et al. 2025a) requires the agent to follow fine-grained navigation instructions to move within a scene. Object Goal Navigation (Chaplot et al. 2020; Wani et al. 2020; Cai et al. 2024; Sadek et al. 2023; Yokoyama et al. 2024) aims to enable agent to efficiently locate specific objects in the environment based on textual descriptions. In contrast, we focus on outdoor scenarios and train the language-guided navigation model using only real-world data, thereby avoiding this gap and enabling the model to learn navigation rules and environmental affordances inherent in urban environments.
@@
Learning from web-scale video data has advanced language and vision tasks (Wang et al. 2024b; He et al. 2024, 2025b). However, a key challenge is the lack of action labels for navigation tasks. LeLaN (Hirose et al. 2024) uses vision-language model prompting and pretrained navigation models to generate action labels, while (Liu et al. 2025b) applies visual odometry to annotate video frames with pose information. However, LeLaN focuses on close-range object navigation within indoor static scenes, making it incapable of handling complex urban environments, and the constructed dataset does not encompass obstacle avoidance capabilities. In contrast, UrbanNav distinguishes itself by implementing a robust filtering pipeline to select robot-compatible data, ensuring viewpoint alignment and safe navigation behaviors. Additionally, our approach generates language-grounded instructions tied to landmarks and trajectories, enabling scalable, cost-effective, and high-quality action label generation for imitation learning in complex urban navigation tasks.
@@ 3.2 Labeling In-the-Wild Videos
We uniformly segment the raw videos into 2-minute clips, each representing a candidate navigation trajectory. To extract pose information, we employ the state-of-the-art visual odometry model DPVO (Teed et al. 2023). The first frame of each clip is taken as the origin of a local world coordinate system, and we annotate the camera pose of each subsequent frame relative to it. Although visual odometry may suffer from cumulative drift across long horizons, this effect is largely mitigated in our case: the policy is trained to predict future actions based on only the past 8 steps, operating within short temporal windows where VO remains locally consistent. This design, following the practice in (Liu et al. 2025b), enables reliable supervision from noisy egocentric videos. As a result, we obtained a total of 106,603 egocentric navigation trajectories with pose labels, spanning 3,553 hours.
@@
A crucial yet often overlooked issue in learning from in-the-wild walking videos is that not all segments are suitable for training embodied agents. Robotic platforms typically employ forward-facing, fixed-view cameras, requiring visual inputs that maintain consistent alignment with motion direction. However, Unlike robots, human walkers exhibit flexible body dynamics and frequently adjust their head orientation, resulting in significant pitch variations and misalignment between the camera viewpoint and the actual trajectory. Such discrepancies render many video clips incompatible with robotic requirements. Additionally, some segments feature densely crowded scenes with close pedestrian interactions, which pose safety risks for robotic deployment. Unlike prior work (Liu et al. 2025b), which often overlooks these issues and risks degrading policy performance, our approach employs a robust filtering pipeline to select predominantly robot-compatible data for training.
First, we estimate the per-frame camera pitch angle and reject trajectories exhibiting excessive vertical oscillations—specifically, those with pitch variation beyond $15^{\circ}$. We further analyze the alignment between movement direction and viewing direction using a sliding window, and discard segments with significant directional divergence (over $60^{\circ}$), which commonly results from abrupt head turns or side glances. Second, to eliminate trajectories captured in overly crowded areas, we apply YOLOv10 (Wang et al. 2024a) to detect pedestrians. Based on empirical observations, we discard any trajectory where more than five people appear in a single frame and such occurrences happen in more than three frames, indicating sustained dense proximity. After applying these filtering steps, we retain 47,008 high-quality, robot-compatible trajectories spanning 1,566 hours, which serve as the core data for training.
@@
We leverage Qwen2.5-VL-72B (Bai et al. 2025), a state-of-the-art VLM, to automatically detect and localize candidate landmarks in video frames via prompted queries aligned with the above criteria. The model outputs both bounding boxes and preliminary landmark names. To ensure quality, we manually review and filter out low-confidence or ambiguous annotations. For the retained landmarks, we prompt Qwen2.5-VL-72B to generate concise and descriptive natural language instructions. Through this process, we obtain a total of 3 million landmark annotations, each paired with a bounding box and a language description. On average, each trajectory contains 65 identified landmarks, with the mean description length being 17 words.
Method Test Seen Test Unseen
@@ 3.3 Policy Architecture and Training
As illustrated in Fig. 3, our policy model predicts a sequence of future egocentric positions based on both language instructions and visual observations. The input to the model comprises four components: (1) the language instruction, (2) the current visual observation, (3) the past $k$ visual observations. It builds upon previous works (Hirose et al. 2024; Liu et al. 2025b). We use CLIP (Radford et al. 2021) to encode the language instruction and DINOv2 (Oquab et al. 2023) to extract features from all visual frames. Both encoders are kept frozen during training. To ground the current observation in the instruction, we apply a FiLM module (Perez et al. 2018) to modulate the current visual embedding using the language features. All input tokens—including language embeddings, FiLM-modulated current visual features and historical visual observation—are concatenated and fed into a Transformer encoder. This encoder captures both temporal dynamics and cross-modal interactions, and outputs a predicted future trajectory in egocentric coordinates. To enable smoother and more anticipatory navigation, we adopt a multi-step prediction framework following prior work such as (Sridhar et al. 2024). At each time step, the model predicts the agent’s waypoints for the next $k=8$ steps using a lightweight action head.
@@ 4.1 Experimental Setup
We compare our model against several prominent policies previously for real-world navigation. To ensure a fair comparison, we adapt these baselines to our language-guided task. For instance, NoMaD (Sridhar et al. 2024) and ViNT (Shah et al. 2023b) were originally designed for image-goal navigation, and lacked native support for textual instructions. To adapt them, we augmented their architectures by encoding the natural language instruction with CLIP (Radford et al. 2021), concatenating the resulting text features with the visual features and passing them through a fusion layer, while the core network remained unchanged. Similarly, LeLaN (Hirose et al. 2024), originally developed for indoor object-goal navigation with textual inputs, predicts robot linear and angular velocities. To align its output space with our task formulation for consistent evaluation, we adapted its output head to directly regress the future navigation waypoints.
@@
Given the challenges of evaluating end-to-end task completion in real-world environments without autoregressive execution, we designed a comprehensive evaluation protocol that includes both offline benchmarking and real-world deployment. (1) Offline Evaluation. For offline evaluation, we assess the model’s performance on a held-out validation set. The model is given a sequence of historical observations, positions, and a natural language instruction. Its task is to predict a future trajectory, and we measure the deviation between this prediction and the ground-truth trajectory. Following prior work (Liu et al. 2025b), we use the average orientation error (AOE) and maximum average orientation error (MAOE) to measure the directional alignment between predicted and ground-truth actions. To better capture local path-following capabilities essential for obstacle avoidance, we introduce two additional metrics: average distance error (ADE), which calculates the mean L2-distance between predicted and ground-truth positions, and maximum average distance error (MADE), which represents the Fréchet distance (Alt and Godau 1995) in a discrete setting. For these metrics, a lower value indicates better performance. A more detailed definition of these metrics is provided in the appendix. (2) Real-World Deployment. To validate the policy’s real-world generalization, we deploy the model on a physical robot platform. The primary metric for this evaluation is the navigation success rate, which measures the percentage of trials where the robot successfully reaches the specified goal landmark without any collisions.